# YUI: Day-ahead Electricity Price Forecasting Using Invariance Simplified Supply and Demand Curve

### v1

@@ Appendix A Datasets Introduction
The assumption of time invariance is closely related to the modeling of thermal power generation. It is important to highlight that our methodology remains applicable in addressing the future’s needs, which include the anticipation of increased deregulation, the integration of renewable resources, and the implementation of energy storage solutions. Due to the high cost of thermal power generation, as long as it exists, it will serve as the marginal price of the supply curve. The instability of renewable energy generation and the inertia of social transition determine that thermal power generation is difficult to be completely eliminated in the foreseeable future, so the time invariance assumption of the supply curve can still apply for a long time. As the proportion of renewable energy generation increases and energy storage technology develops, if thermal power generation has not been completely eliminated at this time, our method still applies according to the previous analysis, and even because of the progress of energy storage technology, the cost of supply during peak and trough periods will be more similar, which actually enhances our assumption of time invariance (the supply curve is more like a horizontal line); if renewable energy occupies all power generation shares, the supply curve at that time may be more affected by storage scheduling costs, which also have time invariance. Of course, these need to be studied in depth after the relevant technologies mature. It is worth mentioning that in the Shanxi market we analyzed, the proportion of renewable energy generation has already reached more than half, far higher than the global average level (30%), and our method still works well. As for the increased market deregulation, our method comes from the principle of supply and demand and is still applicable in a free market. The ISO dataset in the experimental part is heavily deregulated markets, and the electricity price is determined by the transaction price. In summary, we believe that the assumption of time invariance of the supply curve can be established in a wide range of time and space.
Appendix B Tricks Helping Fit Piecewise Linear Function
@@
Appendix C Baseline Details
  • TimesNet (Wu et al., 2022). TimesNet, through its modular structure, decomposes complex time series changes into different periods. By transforming the original one-dimensional time series into a two-dimensional space and using CNN, it unifies the modeling of intra-cycle and inter-cycle changes.
  • Koopa (Liu et al., 2023b). Koopa focuses on describing ubiquitous non-stationary time series. It models time series data from a dynamical perspective and naturally solves the problem of non-linear evolution in real-world time series through modern Koopman theory.
  • Informer (Zhou et al., 2021). Informer, an efficient transformer-based model for Long Sequence Time-series Forecasting (LSTF), addresses the Transformer’s issues of quadratic time complexity, high memory usage, and encoder-decoder architecture limitations. It features a ProbSparse self-attention mechanism for improved time complexity and memory usage, self-attention distilling for handling long input sequences, and a generative style decoder for fast long-sequence predictions.
  • Autoformer (Wu et al., 2021). Autoformer, a novel architecture with an Auto-Correlation mechanism, improves long-term time series forecasting by efficiently discovering dependencies and aggregating representations, outperforming traditional Transformer models and achieving state-of-the-art accuracy across various applications.
  • FEDformer (Zhou et al., 2022). FEDformer is a novel method for time series forecasting that combines the Transformer model with seasonal-trend decomposition and frequency enhancement. This approach not only captures the global trend and detailed structures of time series, but also significantly improves prediction accuracy and efficiency.
  • DLinear (Zeng et al., 2023). DLinear is a simple linear model. According to reports, its performance in the field of time series forecasting can be compared with Transformer-based models.
  • iTransformer (Liu et al., 2023a). iTransformer is a novel Transformer-based architecture for time series forecasting. It considers different variables separately, with each variable being encoded into independent tokens. It uses attention mechanisms to model the correlation between different variables, and feed-forward networks to model the temporal correlation of variables, thereby obtaining a better sequence temporal representation.
  • SARIMA (Zhao et al., 2017; McHugh et al., 2019).The SARIMA (Seasonal Autoregressive Integrated Moving Average) model is a statistical approach used for time series forecasting, which captures autocorrelation, differencing, and seasonality in the data. In day-ahead electricity price prediction, SARIMA can model the time-dependent structure and seasonality of the prices, providing accurate forecasts that are crucial for operational and strategic decisions in the energy market.
  • Linear (Uniejewski et al., 2016; Lago et al., 2021). Linear models are models that assume a linear relationship between the input and output variables. They are widely used in electricity price forecasting.
  • SVM (Che and Wang, 2010; Wang et al., 2017; Prahara and Hariadi, 2022). SVM is a powerful supervised learning algorithm that efficiently perform a non-linear classification using what is called the kernel trick, implicitly mapping their inputs into high-dimensional feature spaces. SVM models are also used in electricity price forecasting.
  • XGBoost (Manfre Jaimes et al., 2023; Xie et al., 2022b).The XGBoost (Extreme Gradient Boosting) model is a machine learning technique that uses gradient boosting framework for regression and classification problems. In day-ahead electricity price prediction, XGBoost can handle non-linear relationships between features and target variable, and it’s robust to outliers, making it a powerful tool for predicting prices with high accuracy.
  • DNN (Yamin et al., 2004; Lago et al., 2018a; Darudi et al., 2015). The DNN is a multi-layer feed forward network that uses a multivariate framework. There are many DNN-based models in electricity price forecasting.
  • LASSO-RF (Ludwig et al., 2015). This model utilizes LASSO for feature selection to enhance the accuracy of electricity price forecasting. A case study for electricity price forecasting is presented, comparing different models. Based on the evaluation of forecasting accuracy, the final model used for price forecasting is Random Forest, which can automatically select important variables and handle non-linear relationships.
  • VMD-LSTM (Xiong and Qing, 2023). This model includes three strategies: an Adaptive Copula-Based Feature Selection (ACBFS) algorithm for input feature selection, a new signal decomposition technique based on a decomposition denoising strategy, and a Long Short-Term Memory (LSTM) model for forecasting.
Appendix D Comparing the Predictability between Price and Load Rate
@@
Figure 6. The ACF plot of the Load Rate and Day-ahead Price variable on Shanxi dataset, where every 96 lags on the x-axis represent one day.
@@
Figure 7. The ACF plot of the Load Rate and Day-ahead Price variable on ISO New England dataset, where every 24 lags on the x-axis represent one day.
@@
Figure 8. Comparison of sMAPE for forecasting Load Rate and Day-ahead Price respectively on the Shanxi and ISO New England datasets using simple forecasting method.
@@
We conducted a comparative analysis of the predictability of two variables, namely Load Rate and Day-ahead Prices, using Autocorrelation Function (ACF) and Symmetric Mean Absolute Percentage Error (sMAPE) derived from a simple forecasting method. The datasets used for this analysis were sourced from Shanxi and ISO New England. It’s important to note that both Load Rate and Day-ahead Prices exhibit daily periodicity. The ACF_daily indicator, which measures predictability, shows that the larger the peak in the ACF curve, the stronger the predictability (Chen et al., 2023). As depicted in Fig. 6(a) and  6(b), the ACF plot for Load Rate in the Shanxi dataset has a larger peak compared to that of the Day-ahead Price, indicating a higher ACF_daily and thus, stronger predictability. This observation is consistent with the results from the ISO New England dataset, as shown in Fig. 7(a) and  7(b). We also employ a simple forecasting method, which involves forecasting the data for a target date using the data from the previous day. This method, effective for variables with daily periodicity, results in a sMAPE error. Fig. 8(a) and  8(b) present the forecast results for Load Rate and Day-ahead electricity Prices using this method. In both datasets, the sMAPE for Load Rate forecasts was smaller. In conclusion, our analysis suggests that Load Rate is a more suitable variable for forecasting.

### v2

@@ CONCLUSION
The assumption of short-term stability is closely related to the modeling of thermal power generation. It is important to highlight that our methodology remains applicable in addressing the future’s needs, which include the anticipation of increased deregulation, the integration of renewable resources, and the implementation of energy storage solutions. Due to the high cost of thermal power generation, as long as it exists, it will serve as the marginal price of the supply curve. The instability of renewable energy generation and the inertia of social transition determine that thermal power generation is difficult to be completely eliminated in the foreseeable future, so the short-term stability assumption of the supply curve can still apply for a long time. As the proportion of renewable energy generation increases and energy storage technology develops, if thermal power generation has not been completely eliminated at this time, our method still applies according to the previous analysis, and even because of the progress of energy storage technology, the cost of supply during peak and trough periods will be more similar, which actually enhances our assumption of short-term stability (the supply curve is more like a horizontal line); if renewable energy occupies all power generation shares, the supply curve at that time may be more affected by storage scheduling costs, which also have short-term stability. Of course, these need to be studied in depth after the relevant technologies mature. It is worth mentioning that in the Shanxi market we analyzed, the proportion of renewable energy generation has already reached more than half, far higher than the global average level (30%), and our method still works well. As for the increased market deregulation, our method comes from the principle of supply and demand and is still applicable in a free market. The ISO dataset in the experimental part is heavily deregulated markets, and the electricity price is determined by the transaction price. In summary, we believe that the assumption of short-term stability of the supply curve can be established in a wide range of time and space.
Appendix C
@@
Appendix D
@@
Figure 9: The ACF plot of the Load Rate and Day-ahead Price variable on Shanxi dataset, where every 96 lags on the x-axis represent one day.
@@
Figure 10: The ACF plot of the Load Rate and Day-ahead Price variable on ISO New England dataset, where every 24 lags on the x-axis represent one day.
@@
Figure 11: Comparison of sMAPE for forecasting Load Rate and Day-ahead Price respectively on the Shanxi and ISO New England datasets using simple forecasting method.
@@
We conducted a comparative analysis of the predictability of two variables, namely Load Rate and Day-ahead Prices, using Autocorrelation Function (ACF) and Symmetric Mean Absolute Percentage Error (sMAPE) derived from a simple forecasting method. The datasets used for this analysis were sourced from Shanxi and ISO New England. It’s important to note that both Load Rate and Day-ahead Prices exhibit daily periodicity. The ACF_daily indicator, which measures predictability, shows that the larger the peak in the ACF curve, the stronger the predictability (Chen et al. 2023a). As depicted in Fig. 9(a) and  9(b), the ACF plot for Load Rate in the Shanxi dataset has a larger peak compared to that of the Day-ahead Price, indicating a higher ACF_daily and thus, stronger predictability. This observation is consistent with the results from the ISO New England dataset, as shown in Fig. 10(a) and  10(b). We also employ a simple forecasting method, which involves forecasting the data for a target date using the data from the previous day. This method, effective for variables with daily periodicity, results in a sMAPE error. Fig. 11(a) and  11(b) present the forecast results for Load Rate and Day-ahead electricity Prices using this method. In both datasets, the sMAPE for Load Rate forecasts was smaller. In conclusion, our analysis suggests that Load Rate is a more suitable variable for forecasting.
Appendix E
In these two datasets, we use the supply and demand quantities data of thermal power generation to construct the supply and demand curve. The day-ahead electricity market comprises several key supplier categories: thermal, wind, solar, nuclear, and hydro power generation (Bichler et al. 2022). The supply curves for these categories exhibit significant variations. The day-ahead electricity market can be approximated as a perfectly competitive market, (OpenStax 2024d, c, b, a) where highest marginal costs equals the price (Hogan 2022). In the electricity market, the highest marginal cost is typically associated with thermal power generation. (Su et al. 2020; Hogan 2022) The ISO New England dataset corroborates this perspective. It documents the supply costs of various energy sources across 108,646 trading intervals from October 1, 2022, to November 11, 2023. The data reveals that in 80.1% of these intervals, the marginal cost is tied to thermal power generation. This suggests a significant role of thermal power in influencing market dynamics.  (iso 2023)This observation aligns well with conventional wisdom for several reasons:
  • Historical precedence. Thermal power generation was established earlier than other forms of energy generation, providing a stable power supply (Yang and Xu 2022; Farnoosh 2022).
  • Market dominance. Thermal power holds a substantial share of the electricity market, reflecting the industry’s maturity (Liu et al. 2021).
  • Resource allocation. The cost of thermal power generation, which relies on higher-cost fossil fuels, exceeds that of photovoltaic, solar, and hydro power generation (Liu, Geng, and Guan 2020).
Appendix F
  • TimesNet (Wu et al. 2022). TimesNet, through its modular structure, decomposes complex time series changes into different periods. By transforming the original one-dimensional time series into a two-dimensional space and using CNN, it unifies the modeling of intra-cycle and inter-cycle changes.
  • Koopa (Liu et al. 2023b). Koopa focuses on describing ubiquitous non-stationary time series. It models time series data from a dynamical perspective and naturally solves the problem of non-linear evolution in real-world time series through modern Koopman theory.
  • Informer (Zhou et al. 2021). Informer, an efficient transformer-based model for Long Sequence Time-series Forecasting (LSTF), addresses the Transformer’s issues of quadratic time complexity, high memory usage, and encoder-decoder architecture limitations. It features a ProbSparse self-attention mechanism for improved time complexity and memory usage, self-attention distilling for handling long input sequences, and a generative style decoder for fast long-sequence predictions.
  • Autoformer (Wu et al. 2021). Autoformer, a novel architecture with an Auto-Correlation mechanism, improves long-term time series forecasting by efficiently discovering dependencies and aggregating representations, outperforming traditional Transformer models and achieving state-of-the-art accuracy across various applications.
  • FEDformer (Zhou et al. 2022). FEDformer is a novel method for time series forecasting that combines the Transformer model with seasonal-trend decomposition and frequency enhancement. This approach not only captures the global trend and detailed structures of time series, but also significantly improves prediction accuracy and efficiency.
  • DLinear (Zeng et al. 2023). DLinear is a simple linear model. According to reports, its performance in the field of time series forecasting can be compared with Transformer-based models.
  • iTransformer (Liu et al. 2023a). iTransformer is a novel Transformer-based architecture for time series forecasting. It considers different variables separately, with each variable being encoded into independent tokens. It uses attention mechanisms to model the correlation between different variables, and feed-forward networks to model the temporal correlation of variables, thereby obtaining a better sequence temporal representation.
  • SARIMA (Zhao et al. 2017; McHugh et al. 2019).The SARIMA (Seasonal Autoregressive Integrated Moving Average) model is a statistical approach used for time series forecasting, which captures autocorrelation, differencing, and seasonality in the data. In day-ahead electricity price prediction, SARIMA can model the time-dependent structure and seasonality of the prices, providing accurate forecasts that are crucial for operational and strategic decisions in the energy market.
  • Linear (Uniejewski, Nowotarski, and Weron 2016; Lago et al. 2021). Linear models are models that assume a linear relationship between the input and output variables. They are widely used in electricity price forecasting.
  • SVM (Che and Wang 2010; Wang et al. 2017; Prahara and Hariadi 2022). SVM is a powerful supervised learning algorithm that efficiently perform a non-linear classification using what is called the kernel trick, implicitly mapping their inputs into high-dimensional feature spaces. SVM models are also used in electricity price forecasting.
  • XGBoost (Manfre Jaimes et al. 2023; Xie et al. 2022).The XGBoost (Extreme Gradient Boosting) model is a machine learning technique that uses gradient boosting framework for regression and classification problems. In day-ahead electricity price prediction, XGBoost can handle non-linear relationships between features and target variable, and it’s robust to outliers, making it a powerful tool for predicting prices with high accuracy.
  • DNN (Yamin, Shahidehpour, and Li 2004; Lago, De Ridder, and De Schutter 2018; Darudi, Bashari, and Javidi 2015). The DNN is a multi-layer feed forward network that uses a multivariate framework. There are many DNN-based models in electricity price forecasting.
  • LASSO-RF (Ludwig, Feuerriegel, and Neumann 2015). This model utilizes LASSO for feature selection to enhance the accuracy of electricity price forecasting. A case study for electricity price forecasting is presented, comparing different models. Based on the evaluation of forecasting accuracy, the final model used for price forecasting is Random Forest, which can automatically select important variables and handle non-linear relationships.
  • VMD-LSTM (Xiong and Qing 2023). This model includes three strategies: an Adaptive Copula-Based Feature Selection (ACBFS) algorithm for input feature selection, a new signal decomposition technique based on a decomposition denoising strategy, and a Long Short-Term Memory (LSTM) model for forecasting.
The detaied experiment hyperparameters can be seen in supplementary material and code.
Appendix G
In the field of electricity price forecasting, the most widely used metrics to measure the accuracy of point forecasts are the Mean Absolute Error (MAE), the Root Mean Square Error (RMSE), and the Mean Absolute Percentage Error (MAPE). In particular, in most electricity trade applications, the underlying risk, profits, and costs depend linearly on the price and on the forecasting errors. Hence, linear metrics represent better than quadratic metrics the underlying risks of forecasting errors (Lago et al. 2021).
The formula for sMAPE is:
$sMAPE=\frac{100\%}{n}\sum_{t=1}^{n}2\frac{|Y_{t}-\hat{Y}_{t}|}{|Y_{t}|+|\hat{Y}_{t}|}$ (2)